# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ 3.1 Phase 1: Annotator Development
Level
All Easy Medium Hard
Type Dataset Attribute
InsightFace
[43]
Face++
[42]
Ours
InsightFace
[43]
Face++
[42]
Ours
InsightFace
[43]
Face++
[42]
Ours
InsightFace
[43]
Face++
[42]
Ours
76.7289 78.0764 83.8978 97.0133 97.0863 99.7333 74.2400 75.356 87.5333 58.9333 61.787 64.4267
Gender (0.4985) (0.4266) (0.3697) (0.1293) (0.3414) (0.1265) (0.8182) (0.5938) (0.5007) (0.5481) (0.3445) (0.4818)
54.4311 58.4889 77.4044 68.000 73.0067 98.4133 49.8000 53.2467 78.9733 45.4933 49.2134 54.8267
Intra- Domain VGGFace2 [47] Age (0.7443) (0.7341) (0.6714) (0.5530) (0.6534) (0.1543) (0.6613) (0.8465) (1.0771) (1.0186) (0.7025) (0.7827)
84.9733 89.1714 91.3000 96.8267 98.0528 98.9333 88.0933 94.3074 98.8667 70.0000 75.1539 76.1000
Gender (0.4651) (0.1974) (0.2058) (0.3832) (0.1483) (0.0943) (0.4668) (0.2586) (0.1033) (0.5452) (0.1854) (0.4197)
59.4867 64.1893 75.5393 71.254 80.5980 93.1980 58.1720 63.8340 81.1960 49.0340 48.1360 52.2240
A-FF++ [19] Age (0.9291) (0.7609) (0.5130) (0.5973) (0.4140) (0.3110) (0.4489) (0.6733) (0.4702) (1.7410) (1.1954) (0.7577)
70.1111 76.0917 78.2922 85.8533 92.1414 96.2933 68.4666 73.5088 76.2005 56.0133 62.6249 62.3334
Gender (0.5037) (0.6290) (0.5178) (0.5239) (0.5447) (0.5927) (0.5667) (0.4910) (0.5028) (0.5014) (0.8513) (0.4580)
66.6967 69.5907 77.1800 72.1580 84.5000 95.3820 64.398 64.238 73.1620 63.5340 60.034 62.9960
A-DFDC [19] Age (0.8015) (0.5687) (0.6300) (0.6785) (0.4908) (0.3247) (0.9182) (0.5423) (0.8592) (0.8078) (0.6730) (0.7061)
66.7156 70.7983 74.9822 85.5467 88.9791 94.0400 60.5467 64.2144 70.0267 54.0533 59.2015 60.8800
A-DFD [19] Gender (0.7681) (1.2229) (0.6029) (0.6791) (0.8297) (0.2999) (0.9017) (1.3436) (0.9471) (0.7235) (1.4953) (0.5616)
91.9244 90.8100 95.1489 98.9733 98.1867 99.9867 94.0000 94.4933 99.7600 82.8000 79.7500 85.7000
Cross- Domain A-Celeb- DF-v2[19] Gender (0.3003) (0.4487) (0.4088) (0.1769) (0.2286) (0.0267) (0.3239) (0.5052) (0.0998) (0.4000) (0.6124) (1.1000)
Table 3: Comparing our annotator against Face++ [42] and InsightFace [43] under intra-domain and cross-domain evaluations (Accuracy (%)) on different levels of sample difficulty. The prediction mean and standard deviation (in parentheses) are reported. The best results are shown in Bold. More results are in Appendix A.1.3.
Evaluation. To demonstrate our annotator’s effectiveness, we will answer the following questions: Q1: How are the general performance and generalization capability of our annotator compared with the baselines? Q2: How does sample difficulty affect the annotator’s performance? In leveraging the good generalization capabilities of CLIP, our annotator is trained on the VGGFace2 [47] dataset, which contains 9.1K individuals with 3.3M images. More importantly,  [48] provides comprehensive demographic annotations for this dataset. We compare our annotator with the current state-of-the-art face attribute prediction tools Face++ [42] and InsightFace [43]. Since they do not offer predictions for the race attribute, our evaluation is confined to gender and age. The mean and standard deviation are reported based on 5 random runs. More details are in Appendix A.1.1.
For Q1, Setting: We perform intra-domain (train on VGGFace2, test on its official test set) and cross-domain (train on VGGFace2, test on four AI-generated face datasets) evaluations. Specifically, A-FF++, A-DFDC, A-DFD, and A-Celeb-DF-v2 are selected from [19] for cross-domain evaluation. Since A-DFD and A-Celeb-DF-v2 have limited age and race annotations, our evaluation of these two is confined to gender. These datasets are chosen because they closely match our objective and are not used to train Face++ and InsightFace. Results: The ‘All’ results in Table 3 demonstrate our annotator’s superiority in general performance and generalization capability against Face++ and InsightFace. Under intra-domain evaluation, it surpasses the second-best method, Face++, by 5.8% on gender and 18.9% on age. In cross-domain evaluation, our annotator maintains high accuracy on all datasets, reflecting good generalization. Remarkably, on the A-FF++ dataset, our annotator outperforms Face++ by a substantial margin of up to 11.4% and InsightFace by 16.1% on age.
For Q2, Setting: We also design a stratified evaluation method by separating each test dataset into three subsets—Easy, Medium, and Hard based on the estimated uncertainty scores. Specifically, for each demographic attribute $a$, we define two thresholds $t_{1}^{a}$ and $t_{2}^{a}$, where $t_{1}^{a}<t_{2}^{a}$ (more details are in Appendix A.1.2). Then, we have $Easy={\{X\mid V(X^{(a)})<t_{1}^{a}\}}$, $Medium=\{X\mid t_{1}^{a}\leq V(X^{(a)})\leq t_{2}^{a}\}$, and $Hard=\{X\mid V(X^{(a)})>t_{2}^{a}\}$. Next, we sample 1,500 images from each subset. This stratification is crucial for a thorough examination of the model’s performance across a broad spectrum of data challenges. To avoid attribute-specific biases, each subset is balanced with respect to attribute. Results: Table 3 illustrates that while all methods show decreased accuracy as the sample difficulty level increases, our annotator demonstrates greater resilience. For example, under intra-domain evaluation, our annotator’s gender performance drops by 10.2% from easy to medium difficulty, compared to Face++’s 21.7% drop. In cross-domain scenario, our annotator experiences a 14.3% reduction on gender in A-Celeb-DF-v2 [19], versus InsightFace’s 16.2% from easy to hard.
3.2 Phase 2: Demographically Annotation Generation
Data Collection. We build our AI-Face dataset by collecting and integrating public AI-generated face images sourced from academic publications, GitHub repositories, and commercial tools. More details are in Appendix A.2.1. Specifically, the fake face images in our dataset originate from 4 Deepfake Video datasets (i.e., A-FF++ [19], A-DFDC [19], A-DFC [19], and A-Celeb-DF-v2 [19]), generated by 10 GAN models (i.e., AttGAN [49], MMDGAN [50], StarGAN [49], StyleGANs [49, 51, 52], MSGGAN [50], ProGAN [53], STGAN [50], and VQGAN [54]), and 8 DM models (i.e., DALLE2 [55], IF [55], Midjourney [55], DCFace [56], Latent Diffusiin [57], Palette [58], Stable Diffusion v1.5 [59], Stable Diffusion Inpainting [59]). This constitutes a total of 1,245,660 fake face images in our dataset. These fake images are correspondingly generated from 8 real source datasets (i.e., FFHQ [6], CASIA-WebFace [37], IMDB-WIKI [38], CelebA [36], and real images from FF++ [2], DFDC [39], DFD [40], and Celeb-DF-v2 [41]). This constitutes a total of 866,096 real face images in our dataset. In general, our dataset contains 30 subsets and 37 generation methods (i.e., 5 in A-FF++, 5 in A-DFD, 8 in A-DFDC, 1 in A-Celeb-DF-v2, 10 GANs, and 8 DMs). We use RetinaFace [60] for detecting and cropping faces to ensure each image only contains one face.
Gender Age Race
Type Datasets ACC(%) Precision(%) Recall(%) ACC(%) Precision(%) Recall(%) ACC(%) Precision(%) Recall(%)
A-FF++ [19] 8.0163 17.3354 5.8314 19.9002 30.6658 29.6071 28.7865 35.7122 41.1687
Ours-FF++ (w/o Correction) 91.9837 82.6646 94.1684 21.1830 32.1232 45.7231 45.9775 50.3803 40.1949
A-DFDC [19] 20.2252 27.5332 21.6538 16.7493 29.0640 29.5519 18.1115 15.1092 22.0637
For Q1 Ours-DFDC (w/o Correction) 79.7748 72.4668 78.3462 45.9748 49.4734 48.7861 70.9001 64.7655 65.1608
Ours (w/o Correction) 83.4167 83.4167 83.4242 43.8333 43.8333 54.1792 67.4167 65.0718 59.2350
For Q2 Ours 84.8333 84.8738 84.8599 44.7500 44.0937 54.6033 68.8333 66.6440 61.3225
For Q3 Ours 98.6667 98.6688 98.6667 56.2500 50.1748 53.0514 86.2500 75.5216 67.4076
Table 4: Evaluation results of our dataset annotation quality for questions Q1, Q2, and Q3. ‘Ours-FF++ (w/o Correction),’ ‘Ours-DFDC (w/o Correction),’ and ‘Ours (w/o Correction)’ represent our predicted annotations on A-FF++, A-DFDC, and our entire dataset without human correction, respectively. ACC represents Accuracy.
Annotator Prediction. For our collected images, annotation generation is iterative, integrating uncertainty scores into each prediction by our annotator in Phase 1, as shown in Fig. 2.
Human Correction. As described in ‘Uncertainty Estimation’ in Section 3.1, the annotator may mispredict ambiguous face images, necessitating human review and correction. To this end, we propose two annotation correction strategies: 1) For subsets that have the same images and demographic attribute classes as those in existing datasets, such as A-FF++ [19] and A-DFDC [19], we filter out images that may need human correction based on annotation inconsistency.
Figure 3: Uncertainty-based strategy for identifying ambiguous faces for human correction.
2) For the rest of the subsets, we identify the most ambiguous images that need human correction based on uncertainty scores. Specifically, for demographic attribute $a_{j}$ on subset $j$, we define a specific threshold $t^{a_{j}}$ (more details are in Appendix A.2.2). If $V(X^{(a_{j})})>t^{a_{j}}$, the annotation for attribute $a_{j}$ of the image $X$ will undergo a verification process, potentially requiring human re-annotation (see Fig. 3). In practice, we recruit three humans to correct the filtered images, consolidating their evaluations with a majority vote to finalize annotations.
Evaluation. To estimate our dataset’s quality, we will answer the following questions: Q1: Can we directly incorporate the existing annotations into our dataset? Q2: How is the effectiveness of human correction? Q3: How is the overall annotations’ quality of our dataset?
For Q1, Setting: We compare our dataset’s annotation quality before human correction on A-FF++ (i.e., Ours-FF++ (w/o Correction)) and A-DFDC (i.e., Ours-DFDC (w/o Correction)) against their existing annotation from [19]. We regard human re-labeled annotations as the ground truth. Results: The results in Table 4 ‘For Q1’ show superior annotation accuracy of our datasets. For example, Ours-FF++ (w/o Correction) surpasses A-FF++ by 83.97% in gender accuracy, and Ours-DFDC (w/o Correction) exceeds A-DFDC by 59.55%. The large performance indicates that identified images by annotation inconsistency are mislabeled in A-FF++ [19] and A-DFDC [19], and thus cannot be directly merged into our dataset. Some examples are shown in Appendix A.2.3.
For Q2, Setting: We consider two dataset versions: 1) Ours (w/o Correction), where annotations are not corrected by humans. 2) Ours, where annotations are corrected by humans. With the help of the uncertainty score, we sample 1,200 attribute-balanced images (400 easy, 400 medium, and 400 hard) from the whole dataset to ensure a fair evaluation. Three humans re-annotated these images to establish ground truth. Results: Table 4 ‘For Q2’ shows that human corrections improve performance across all attributes, increasing accuracy by 1.42% for gender, 0.92% for age, and 1.42% for race, validating the effectiveness of our correction strategy. More results see Appendix A.2.4.
For Q3, Setting: We randomly sample 1,200 images from the whole dataset. Three humans also re-annotated these images to create ground truth. Results: As shown in Table 4 ‘For Q3’, Ours reflects the approximate overall annotation quality of our dataset. Notably, the annotations of gender and race attributes show high correctness (e.g., 98.6667% ACC on gender and 86.2500% ACC on race). However, the age annotation shows a lower accuracy since it is challenging to differentiate.
4 Fairness Benchmark Experiments
In this section, we estimate the existing AI-generated image detectors’ fairness performance alongside their utility on our AI-Face Dataset (80%/20% for Train/Test). Our goal is to show the significance of our dataset and expose the fairness issues of recent detectors in combating AI-generated faces.
Detection Methods. Our benchmark has implemented 12 detectors, as detailed in Appendix B.1. The methodologies cover a spectrum that is specifically tailored to detect AI-generated faces from Deepfake Videos, GANs, and DMs. They can be classified into four types: Naive detectors: refer to backbone models that can be directly utilized as the detector for binary classification, including CNN-based (i.e., Xception [61] and EfficientB4 [62]) and transformer-based (i.e., ViT-B/16 [63]). Frequency-based: explore the frequency domain for forgery detection (i.e., F3Net [64], SPSL [65], and SRM [66]). Spatial-based: focus on mining spatial characteristics (e.g., texture) within images for detection (i.e., UCF [16], UnivFD [67], and CORE [68]). Fairness-enhanced: focus on improving fairness in AI-generated face detection by designing specific algorithms (i.e., DAW-FDD [20], DAG-FDD [20], and PG-FDD [21]). Implementation and training details refer to Appendix B.2.
@@ 4 Fairness Benchmark Experiments
$F_{MEO}(\%)$ 0.387 1.176 0.187 0.279 0.454 0.533 0.305 0.458 1.635 0.404 0.272 0.236
$F_{DP}(\%)$ 2.843 2.052 2.489 2.941 2.998 2.433 2.890 2.456 1.977 2.979 2.799 2.614
$F_{OAE}(\%)$ 0.271 0.595 0.422 0.086 0.188 0.268 0.169 0.557 0.977 0.123 0.192 0.134
Gender $F_{EO}(\%)$ 0.439 1.229 0.235 0.552 0.577 0.536 0.346 0.490 1.846 0.699 0.407 0.237
$F_{MEO}(\%)$ 4.386 8.307 13.078 3.098 4.736 5.470 3.188 14.663 16.001 3.461 3.344 1.956
$F_{DP}(\%)$ 18.248 19.691 18.446 18.282 18.822 16.182 18.770 23.542 24.163 18.306 18.288 18.040
$F_{OAE}(\%)$ 3.509 5.659 5.351 2.217 2.201 4.044 1.847 6.505 5.105 1.365 1.847 1.132
Race $F_{EO}(\%)$ 10.863 19.921 24.002 7.052 7.282 11.602 6.311 30.947 24.015 6.948 6.439 4.039
$F_{MEO}(\%)$ 1.695 3.028 8.931 1.319 1.025 1.090 0.854 5.818 6.964 2.838 0.809 0.781
$F_{DP}(\%)$ 6.242 6.724 6.264 6.357 6.340 5.905 6.257 6.260 5.030 6.249 6.140 6.098
$F_{OAE}(\%)$ 1.028 1.619 3.948 1.017 0.710 0.934 0.635 4.966 3.652 2.610 0.606 0.506
Age $F_{EO}(\%)$ 4.116 6.080 12.888 3.696 2.827 3.116 2.479 15.252 8.382 7.361 2.171 1.587
$F_{MEO}(\%)$ 7.113 9.999 14.667 4.739 7.320 9.731 4.606 17.606 19.303 5.316 4.708 2.604
$F_{DP}(\%)$ 20.675 20.963 20.114 20.492 20.242 19.112 20.704 24.366 25.892 20.373 19.940 20.402
$F_{OAE}(\%)$ 6.174 9.181 7.711 3.692 3.744 6.498 3.061 11.802 6.035 2.641 3.174 1.830
Intersection $F_{EO}(\%)$ 24.520 42.330 49.075 16.699 16.257 25.983 13.932 68.449 47.016 14.539 14.118 8.618
Individual $F_{IND}(\%)$ 112.067 585.935 0.125 46.083 22.982 1.383 3.246 8.606 0.598 28.437 13.706 0.477
Avg-$F_{R}$ 6.824 9.529 7.706 5.941 6.647 5.471 4.353 9.941 9.235 6.118 3.765 2.059
Fairness↓ Avg-$F_{MR}$ 8.020 6.020 7.843 3.981
ACC(%) 97.639 95.404 93.719 98.229 98.274 97.978 98.635 90.229 96.087 97.316 98.543 99.079
AUC(%) 99.768 99.117 98.914 99.826 99.786 99.767 99.885 96.030 98.846 99.703 99.871 99.937
AP(%) 99.846 99.359 99.240 99.885 99.853 99.829 99.917 96.973 98.987 99.802 99.916 99.956
Utility↑ - EER(%) 2.388 4.794 5.829 1.741 1.610 2.134 1.365 10.680 4.656 2.701 1.365 1.212
Training Time / Epoch 1h35min 3h07min 3h26min 1h41min 1h37min 4h05min 5h10min 5h07min 1h36min 1h45min 1h38min 7h45min
Table 5: Overall performance. Top 3 values on each metric are highlighted in green, blue, and yellow.
@@
Evaluation Metrics. To provide a comprehensive benchmarking, we consider 5 fairness metrics commonly used in fairness community [69, 70, 71, 72, 73] and 4 widely used utility metrics. For fairness metrics, we consider Demographic Parity ($F_{DP}$[69, 70], Max Equalized Odds ($F_{MEO}$[72], Equal Odds ($F_{EO}$[71], and Overall Accuracy Equality ($F_{OAE}$[72] for evaluating group (e.g., gender) and intersectional (e.g., individuals of a specific race and simultaneously a specific gender) fairness. We also use individual fairness ($F_{IND}$[73, 74] (i.e., similar individuals should have similar predicted outcomes) for estimation. Fairness metrics definition can be found in Appendix B.3. To compare detectors’ performance clearly and fairly, we define the Average Fairness Rank (Avg-$F_{R}$), which ranks each detector on each fairness metric and averages these ranks. We also define Avg-$F_{MR}$ for the average rank across methods within a model type. For utility metrics, we employ Accuracy (ACC), the Area Under the ROC Curve (AUC), Average Precision (AP), and Equal Error Rate (EER).
Results. Overall Performance. Table 5 reports the overall performance on our AI-Face test set. Our observations are: 1) Most detectors do not have fairness except for Fairness-enhanced detectors, which demonstrate relatively lower performance disparities. 2) The top 3 performing methods are PG-FDD [21], DAG-FDD [20], and UCF [16] according to Avg-$F_{R}$. 3) According to Avg-$F_{MR}$, Fairness-enhanced detectors demonstrate superior performance. Frequency detectors surpass both Spatial and Naive detectors. A possible reason is that frequency features are more focused on the forgery trace while weakening the demographic features. This highlights a potential avenue for future research to enhance detector fairness by integrating frequency features with fairness-enhanced algorithms. 4) 9 out of 12 detectors have an AUC higher than 99%, demonstrating our AI-Face dataset is significant for training AI-face detectors in resulting high utility. 5) PG-FDD demonstrates superior performance but has a long training time, which can be explored and addressed in the future.
Performance on Different Subsets. Fig. 4 demonstrates the intersectional $F_{EO}$ and AUC performance of detectors on each test subset (e.g., subsets originate from different generative methods). We observe that the fairness performance varies a lot among different generative methods in every detector. The largest bias on most detectors comes from detecting face images generated by STGAN [75] and Commercial Tools (CT), including DALLE2 [55], IF [55], and Midjourney [55]. Moreover, the stable utility demonstrates our dataset’s expansiveness and diversity, enabling effective training to detect AI-generated faces from various generative methods. Full evaluation results are in Appendix B.4.
Performance on Different Subgroups. We conduct an analysis of all detectors on intersectional subgroups: Male-White (M-W), Male-Black (M-B), Male-Asian (M-A), Male-Others (M-A), Female-White (F-W), Female-Black (F-B), Female-Asian (F-A), Female-Others (F-O). As shown in Fig. 5, it plots the ratios of FPR for each subgroup to a reference group (M-W). 1) It is clear that facial images of M-A, F-B, and F-A are more likely to be mistakenly detected as fake than facial images of M-W. 2) However, the FPR of M-W is higher than others in DAW-FDD. This highlights a challenge in algorithmic fairness methods: improving performance for minority groups can inadvertently raise the error rate for the majority group (e.g., M-W). See demographic distribution in Appendix A.2.1.

### v3

@@ 3.2.2 Gender and Age Annotation Generation
Annotator Architecture. We build a lightweight annotator based on the CLIP [37] foundation model by leveraging its strong zero-shot and few-shot learning capabilities. Specifically, our annotator employs a frozen pre-trained CLIP ViT L/14 [73] as a feature extractor $\mathbf{E}$ followed by a trainable classifier parameterized by $\theta$, which contains 3-layer Multilayer Perceptron (MLP) $\mathbf{M}$ and a classification head $h$.
Learning Objective. Aware that neural networks can perform poorly when the training dataset suffers from class-imbalance [74] and CLIP is not free from demographic bias [75, 76, 77], we introduce an imbalance loss and fairness loss to address these challenges in the annotator training. Specifically, for image $X_{i}$, its feature $f_{i}$ is obtained through $f_{i}\!\!=\!\!\mathbf{M}(\mathbf{E}(X_{i}))$. Next, two losses are detailed below.
Imbalance Loss: To mitigate the impact of imbalance data, we use Vector Scaling [78] loss, which is a re-weighting method for training models on the imbalanced data with distribution shifts and can be expressed as
$L_{imb}=\frac{1}{n}\sum_{i=1}^{n}-u_{A_{i}}\log\frac{e^{\zeta_{A_{i}}h(f_{i})_{A_{i}}+\Delta_{A_{i}}}}{\sum_{A\in\mathcal{A}}e^{\zeta_{A}h(f_{i})_{A}+\Delta_{A}}},\vspace{-1mm}$
where $u_{A_{i}}$ is the weighting factor for attribute $A_{i}$. $h(f_{i})_{A_{i}}$ is the predict logit on $A_{i}$. $\zeta_{A_{i}}$ is the multiplicative logit scaling factor, calculated as the inverse of $A_{i}$’s frequency. $\Delta_{A_{i}}$ is the additive logit scaling factor, calculated as the log of $A_{i}$ probabilities. More details about them are in appendix B.4.
Fairness Loss: We introduce a fairness loss to minimize the disparity between the distribution $\mathcal{D}^{f}$ of $f$ and the conditional distribution $\mathcal{D}^{f_{A}}$ of $f$ on attribute $A\in\mathcal{A}$. Specifically, we follow [79, 80] to minimize the summation of the following Sinkhorn distance between these two distributions:
$L_{fair}=\sum_{A\in\mathcal{A}}\inf_{\gamma\in\Gamma(\mathcal{D}^{f},\mathcal{D}^{f_{A}})}\big{\{}\mathbb{E}_{X\sim\gamma}[c(p,q)]+\alpha H(\gamma|\mu\otimes\nu)\big{\}},$
where $\Gamma(\mathcal{D}^{f},\mathcal{D}^{f_{A}})$ is the set of joint distributions based on $\mathcal{D}^{f}$ and $\mathcal{D}^{f_{A}}$. Let $p$ and $q$ be the points from $\mathcal{D}^{f}$ and $\mathcal{D}^{f_{A}}$, respectively. Then, $c(p,q)$ represents the transport cost [80]. Let $\mu$ and $\nu$ be the reference measures from the set of measures on $f$. Then, $H(\gamma|\mu\otimes\nu)$ represents the relative entropy of $\gamma$ with respect to the product measure $\mu\otimes\nu$. $\alpha\geq 0$ is a regularization hyperparameter. In practice, we use the empirical form of $L_{fair}$.
Total Loss: Therefore, the final learning objective becomes $\mathcal{L}(\theta)=L_{imb}+\lambda L_{fair}$, where $\lambda$ is a hyperparameter.
Figure 3: Distribution of face images of the AI-Face dataset. The figure shows the (a) subset distribution and the demographic distribution for (b) skin tone, (c) gender, and (d) gender. The outer rings in (b), (c), and (d) represent the proportion of groups within each attribute category, while the inner rings indicate the distribution of fake (F) and real (R) images within those groups.
Train. Traditional optimization methods like stochastic gradient descent can lead to poor model generalization due to sharp loss landscapes with multiple local and global minima. To address this, we use Sharpness-Aware Minimization (SAM) [81] to enhance our annotator’s generalization by flattening the loss landscape. Specifically, flattening is attained by determining the optimal $\epsilon^{*}$ for perturbing model parameters $\theta$ to maximize the loss, formulated as: $\epsilon^{*}=\arg\max_{\|\epsilon\|_{2}\leq\beta}{\mathcal{L}}\textbf{(}\theta+\epsilon\textbf{)}\approx\arg\max_{\|\epsilon\|_{2}\leq\beta}\epsilon^{\top}\nabla_{\theta}\mathcal{L}=\beta\texttt{sign}(\nabla_{\theta}\mathcal{L})$, where $\beta$ controls the perturbation magnitude. The approximation is based on the first-order Taylor expansion with assuming $\epsilon$ is small. The final equation is obtained by solving a dual norm problem, where sign represents a sign function and $\nabla_{\theta}\mathcal{L}$ being the gradient of $\mathcal{L}$ with respect to $\theta$. As a result, the model parameters are updated by solving: $\min_{\theta}\mathcal{L}\textbf{(}\theta+\epsilon^{*}\textbf{)}$.
Inference. We use the trained annotators to predict demographic labels for each image in AI-Face dataset, except for those from IMDB-WIKI, which already contain true labels.
3.3 Dataset Statistics
Fig. 3 illustrates the subset distribution and demographic attributes of the AI-Face dataset. The dataset contains approximately three times more generated images than real images, with diffusion model-generated images constituting the majority. In terms of demographic attributes, the majorities in skin tone are Tone 5 (31.14%) and Tone 6 (35.16%). The lightest skin tones (Tones 1-3) are underrepresented, comprising only 0.97% of the dataset. The dataset is relatively balanced across gender. Adult (25-44) (49.67%) is the predominant representation in age groups.
3.4 Annotation Quality Assessment
To assess the quality of demographic annotations in our AI-Face dataset, we conducted a user study. Three participants label the demographic attributes for the given images (the details of labeling activities are in appendix B.5), with the final ground truth determined by majority vote. We then compare our annotations with those in A-FF++, A-DFDC, A-CelebDF-V2, and A-DFD datasets. Specifically, we perform two assessments: 1) Strategic comparison: We select 1,000 images from A-FF++ and A-DFDC that have different annotations from AI-Face. These images likely represent challenging cases. 2) Random comparison: We randomly sampled 1,000 images from A-Celeb-DF-V2 and A-DFD. Due to the limited age classes in these datasets, only gender was evaluated. The results, presented in Table 3, demonstrate the high correctness of the AI-Face annotations and their superior quality compared to the annotations of other datasets. For example, our annotation quality (ACC) surpasses those in A-FF++ by 78.714% on gender and 48.000% on age.
4 Fairness Benchmark Settings
This section demonstrates the fairness benchmark settings for detection methods and evaluation metrics on AI-Face (80%/20%: Train/Test). More settings are in Appendix C.1.
Detection Methods. Our benchmark has implemented 12 detectors. The methodologies cover a spectrum that is specifically tailored to detect AI-generated faces from Deepfake Videos, GANs, and DMs. They can be classified into four types: Naive detectors: refer to backbone models that can be directly utilized as the detector for binary classification, including CNN-based (i.e., Xception [82], EfficientB4 [83]) and transformer-based (i.e., ViT-B/16 [84]). Frequency-based: explore the frequency domain for forgery detection (i.e., F3Net [85], SPSL [86], SRM [87]). Spatial-based: focus on mining spatial characteristics (e.g., texture) within images for detection (i.e., UCF [26], UnivFD [88], CORE [89]). Fairness-enhanced: focus on improving fairness in AI-generated face detection by designing specific algorithms (i.e., DAW-FDD [29], DAG-FDD [29], PG-FDD [30]).
  Gender Age
Evaluation Type Dataset ACC Precision Recall ACC Precision Recall
A-FF++ 8.143 17.583 5.966 37.700 39.459 45.381
AI-Face 86.857 74.404 77.367 85.700 74.024 63.751
A-DFDC 21.600 28.604 23.082 33.400 38.011 40.165
Strategic AI-Face 91.700 92.129 83.448 77.000 76.184 62.646
A-Celeb-DF-V2 89.628 90.626 90.494 -
AI-Face 91.206 91.474 91.767 -
A-DFD 70.900 71.686 74.435 -
Random AI-Face 92.300 91.060 91.727 -
Table 3: Annotation quality assessment results (%) for A-FF++, A-DFDC, A-Celeb-DF-V2, A-DFD, and our AI-Face. ACC: Accuracy.
@@ 4 Fairness Benchmark Settings
$F_{MEO}$ 8.836 8.300 6.264 19.938 8.055 10.002 17.325 2.577 10.779 14.118 6.551 6.465
$F_{DP}$ 9.751 6.184 7.728 12.876 9.379 10.897 12.581 8.556 10.317 10.706 8.617 9.746
$F_{OAE}$ 1.271 4.377 2.168 2.818 1.135 0.915 1.883 2.748 1.332 1.667 1.388 0.882
Skin Tone $F_{EO}$ 12.132 11.062 8.813 23.708 9.789 14.239 21.92 5.536 13.069 16.604 7.383 9.115
$F_{MEO}$ 3.975 5.385 5.104 4.717 4.411 6.271 5.074 4.503 5.795 5.510 5.910 3.190
$F_{DP}$ 1.691 1.725 1.344 1.864 1.827 1.957 1.736 1.190 2.154 2.015 2.151 1.252
$F_{OAE}$ 0.975 1.487 1.803 1.129 1.037 1.772 1.451 1.622 1.389 1.325 1.420 1.071
Gender $F_{EO}$ 4.143 5.863 6.031 4.870 4.534 6.78 5.510 5.408 5.931 5.696 6.066 3.702
$F_{MEO}$ 27.883 6.796 14.937 38.801 27.614 24.843 47.500 5.436 33.882 45.466 15.229 14.804
$F_{DP}$ 10.905 11.849 11.839 14.906 11.232 11.570 17.049 15.249 12.564 14.106 9.633 10.467
$F_{OAE}$ 7.265 2.856 6.838 10.116 7.270 6.524 11.652 3.793 8.760 11.878 5.533 5.009
Age $F_{EO}$ 42.216 10.300 30.795 55.032 40.943 38.528 67.545 14.148 48.729 64.384 30.182 29.585
$F_{MEO}$ 10.505 17.586 9.384 21.369 10.379 15.142 20.134 6.119 15.34 16.565 12.178 9.578
$F_{DP}$ 14.511 8.607 11.535 17.175 13.259 15.186 17.03 14.026 14.301 14.088 11.705 14.697
$F_{OAE}$ 2.536 8.461 4.928 4.870 2.464 3.998 3.536 6.287 2.775 3.547 4.035 3.062
Intersection $F_{EO}$ 24.315 25.114 27.443 47.783 21.679 30.112 43.376 20.255 28.84 33.122 26.295 18.348
Fairness(%)↓ Individual $F_{IND}$ 10.338 25.742 0.022 1.872 2.518 7.621 0.767 3.523 0.041 3.772 0.901 0.780
AUC↑ 98.583 98.611 98.69 98.714 98.747 97.936 98.082 98.192 98.579 97.811 98.771 99.172
ACC↑ 96.308 94.203 94.472 95.719 96.346 95.092 95.151 93.651 96.224 95.426 95.722 96.174
AP↑ 99.350 99.542 99.571 99.453 99.356 99.172 99.273 99.400 99.360 99.015 99.498 99.694
EER↓ 5.149 6.689 6.372 5.256 4.371 6.483 7.708 7.633 5.145 7.063 5.499 4.961
Utility(%) - FPR↓ 12.961 20.066 16.426 14.679 13.661 15.746 13.646 18.550 13.410 16.670 14.844 10.971
Training Time / Epoch 1h15min 2h25min 2h40min 1h18min 1h20min 3h10min 5h05min 4h 1h16min 1h25min 1h17min 7h20min
Table 4: Overall performance comparison of difference methods on the AI-Face dataset. The best performance is shown in bold.
@@
Evaluation Metrics. To provide a comprehensive benchmarking, we consider 5 fairness metrics commonly used in fairness community [90, 91, 92, 93, 94] and 5 widely used utility metrics. For fairness metrics, we consider Demographic Parity ($F_{DP}$[90, 91], Max Equalized Odds ($F_{MEO}$[93], Equal Odds ($F_{EO}$[92], and Overall Accuracy Equality ($F_{OAE}$[93] for evaluating group (e.g., gender) and intersectional (e.g., individuals of a specific gender and simultaneously a specific skin tone) fairness. In experiments, the intersectional groups are Female-Light (F-L), Female-Medium (F-M), Female-Dark (Dark), Male-Light (M-L), Male-Medium (M-M), and Male-Dark (M-D), where we group 10 categories of skin tones into Light (Tone 1-3), Medium (Tone 4-6), and Dark (Tone 7-10) for simplicity according to [95]. We also use individual fairness ($F_{IND}$[94, 96] (i.e., similar individuals should have similar predicted outcomes) for estimation. For utility metrics, we employ the Area Under the ROC Curve (AUC), Accuracy (ACC), Average Precision (AP), Equal Error Rate (EER), and False Positive Rate (FPR).
5 Results and Analysis
In this section, we estimate the existing AI-generated image detectors’ fairness performance alongside their utility on our AI-Face Dataset. More results can be found in Appendix D.
5.1 General Fairness Comparison
Overall Performance. Table 4 reports the overall performance on our AI-Face test set. Our observations are: 1) Fairness-Enhanced Models (specifically PG-FDD [30]) are the most effective in achieving both high fairness and utility, underscoring the effectiveness of specialized fairness-enhancement techniques in mitigating demographic biases. 2) UnivFD [88], based on the CLIP backbone [73], also achieves commendable fairness, suggesting that foundation models equipped with fairness-focused enhancements could be a promising direction for developing fairer detectors. 3) Naive detectors, such as EfficientB4 [83], trained on large, diverse datasets (e.g., our AI-Face) can achieve competitive fairness and utility, highlighting the potential of fairness improvements by choosing specific architecture. 4) 10 out of 12 detectors have an AUC higher than 98%, demonstrating our AI-Face dataset is significant for training AI-face detectors in resulting high utility. 5) PG-FDD demonstrates superior performance but has a long training time, which can be explored and addressed in the future.